What is the separation energy for electron present in first excited state of Be+3
 

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Separation energy =13.6eV\times \dfrac{Z^2}{n^2}
Therefore that for triply charged Be atom is13.6eV\times \dfrac{4^2}{2^2}=54.4eV

Regards.